OpenStudy (anonymous):

4(-2)² + 8(-2) + 3(-2) + 6

OpenStudy (anonymous):

=4(4) + -16+-6 +6 =16-16 =0

OpenStudy (anonymous):

Im great at math ! But cut to the chase deeeezzzz is right The answer is 0